# captionUnknown document class

### v2

@@ III-C Hardware Results
IV Analysis of Fall Prediction Framework Performance
@@
Given that the fall prediction algorithm was trained on abrupt and incipient faults introduced by applying pushing forces to region F at the torso back, as shown in Figure 5, and with features defined exclusively within the sagittal plane, it is not expected to perform well on faults outside its training distribution, such as those introduced by pushing the robot on its sides [32]. However, since the robot training data also included several trajectories where the robot rocked before falling on its front or back along the sagittal plane, it raises the question of whether the algorithm can handle faults introduced by applying forces at different locations and angles, such as the front of the torso. To address this, we evaluate the performance of the three components of the fall prediction algorithm across a broad range of faults. We limit our analysis here to abrupt and incipient faults for reasons similar to those discussed in Section III. Additionally, the evaluation is conducted offline. However, given the negligible difference in performance between online and offline execution, as presented in Section III, we can reasonably conclude that the offline results presented here are indicative of the algorithm’s performance in real-time, online scenarios. Furthermore, the results in [1] suggest that the performance of the individual components directly translates to the overall performance of the algorithm. Note that the Digit open-source dataset, [28], will be extended to include the trajectories used here.
IV-A Data Generation
@@
To assess the algorithm’s robustness across a wider range of faults, abrupt and incipient faults are emulated using pushing forces as described in [1]. The test dataset is generated by applying both angled and non-angled forces at 16 distinct locations on the robot, distributed across the sagittal and frontal planes. These locations are grouped into four regions: torso back, torso front, right side, and left side, with five points on the back, five on the front, and three on each side. Figure 5 illustrates the regions and points where the forces are applied.
@@
For each of the 16 locations, 100 trajectories are generated for both abrupt and incipient faults, resulting in 200 trajectories per location, using both angled and non-angled pushes. For convenience, we term critical faults that are introduced by angled and non-angled forces as critical angled and non-angled faults, respectively. Angled forces are applied by selecting the magnitudes of the x and y components of the force, as well as the angle of the force relative to the transverse plane of the robot. The x-component magnitude is chosen such that applying a non-angled x force causes a fall in approximately half of the trajectories (for both the torso back and front). Similarly, the y-component magnitude is set to produce a fall in about half of the trajectories when non-angled y forces are applied (for the torso sides). The angle of the force relative to the transverse plane is randomly selected from a uniform distribution within the range of [-45°, 45°]. These angles were determined through empirical testing, where the physical robot was pushed at various angles to identify the range within which it could be effectively disturbed. Similar to [1], to ensure slight oscillations are present in the robot’s standing posture, random impulsive forces are introduced, but only windows during and after the fault introduction are used for fall prediction.
@@ IV-B Nominal Critical Fault Classifier Results
In [1], the critical fault classifier demonstrated impressive results by detecting abrupt and incipient faults, achieving an average lead time of 1.52 seconds, an average response time of 0.44 seconds, and an average false positive rate of 0 when evaluated on unsafe trajectories derived in a similar manner to its training set. We refer to these results as in-distribution results.
When assessed on a broader range of abrupt and incipient critical faults, as described in the previous section, the classifier achieves an average response time comparable to the in-distribution results, as shown in Table II. While its average lead times decrease by 0.17s and 0.09s for non-angled and angled critical faults, respectively, the corresponding average false positive rates also rise to 0.08 for both cases. This indicates a decline in the nominal critical fault classifier’s performance. This decline in performance is particularly apparent for side pushes, as the false positive rates for critical faults applied to the sides of the robot are on average 0.13 for both angled and non-angled critical faults, whereas for torso-based critical faults, the rates are 0.02 and 0.04 for non-angled and angled forces, respectively. These results align with expectations, as the model’s training data primarily focuses on faults introduced along the sagittal plane and does not adequately capture the variability of side or angled faults.
@@
Metrics
Non
Ang
Ang.
In
Dist
Non
Ang
Ang
In
Dist
FPR
0.08 0.08 0.0 0.0 0.003 0.0
Lead
Time
1.35 1.43 1.52 1.18 1.26 1.18
Response
Time
0.42 0.41 0.44 0.59 0.59 0.62
TABLE II: Nominal [1] and Fine-Tuned Critical Fault Classifier results evaluated on non-angled and angled abrupt and incipient faults introduced by forces applied at various locations on the robot. Acronyms used: Non ang (non-angled), ang (angled), and in dist (in-distribution or tests from [1]).
@@ IV-C Nominal Lead time classifier
When evaluated on in-distribution critical faults, the lead time classifier achieved accuracies of 1.0, 0.95, and 0.74 for the $[0,1]$, $(1,2]$, and $(2,H]$ lead time intervals, respectively [1]. However, similar to the critical fault classifier, the lead time classifier performs worse when evaluated on a broader range of critical faults. For non-angled critical faults, the classifier, as shown in Table III, achieves accuracies of 0.82, 0.57, and 0.61 for the $[0,1]$, $(1,2]$, and $(2,H]$ intervals, respectively. For angled critical faults, the classifier achieves accuracies of 0.84, 0.63, and 0.55 for the same intervals. This decline in performance, as shown in Figure 6, is particularly pronounced for side critical faults. These results mirror those of the critical fault classifier and are consistent with expectations, given that the training data consists solely of faults introduced along the sagittal plane. Furthermore, the lead time classifier achieves its highest accuracy within the $[0,1]$ interval, which is consistent with the findings in [1], where it is noted that the number of data points decreases exponentially as lead time increases. The asymmetry observed in Figure 6 between the left and right sides can be attributed to randomness in data generation, leading to varying force magnitudes even within the same range. When critical faults from one side are emulated on the other, the lead time classifier produces similar results.
@@
Intervals
Non
Ang
Ang
In
Dist
Non
Ang
Ang
In
Dist
$[0,1]$
0.82 0.84 1.0 0.96 0.95 0.98
$(1,2]$
0.57 0.63 0.95 0.84 0.80 0.82
$[2,H)$
0.61 0.55 0.74 0.30 0.27 0.44
TABLE III: Nominal [1] and Fine-Tuned Lead Time Classifier results evaluated on non-angled and angled abrupt and incipient faults introduced by forces applied at various locations on the robot. Acronyms used: Non ang (non-angled), ang (angled), and in dist (in-distribution or tests from [1]).
@@ IV-D Nominal Lead time regressor
When evaluated on in-distribution critical faults within the $[0,1]$ lead time range, the lead time regressor shows minimal error, with a maximum difference of 0.09s between predicted and actual lead time, and a mean and median difference of just 0.01s [1]. However, similar to both the critical fault classifier and the lead time classifier, the regressor performs significantly worse on non-angled and angled critical faults. For non-angled critical faults, the maximum lead time prediction error increases to 1.79s, while the mean and median errors rise to 0.23s and 0.07s, respectively, as shown in Table IV. For critical angled faults, the maximum error increases to 2.41s, with mean and median errors of 0.25s and 0.11s, respectively. Given that the regressor’s objective is to predict lead time within the $[0,1]$ interval, these error increases are unacceptable. Despite these significant errors, as shown in Figure 7, the regressor performs better for the torso back non-angled critical faults compared to the other regions and faults.
@@
Figure 7: Regressor results when evaluated on non-angled and angled critical faults across various locations. The circles and squares indicate the non-angled and angled critical faults, respectively.
@@
Differences
Non
Ang
Ang
In
Dist
Non
Ang
Ang
In
Dist
Max
1.79 2.41 0.09 0.23 0.32 0.18
Mean
0.23 0.25 0.01 0.03 0.03 0.04
Median
0.07 0.11 0.01 0.02 0.02 0.03
TABLE IV: Nominal and Fine-Tuned Lead Time Regressor results evaluated on non-angled and angled abrupt and incipient faults introduced by forces applied at various locations on the robot. Acronyms used: Non ang (non-angled), ang (angled), and in dist (in-distribution or tests from [1]).
V Fine-Tuning the Nominal Fall Prediction Algorithm
This section explores strategies to enhance the robustness of the nominal fall prediction algorithm against omnidirectional faults, addressing its suboptimal performance, particularly when faults are introduced by forces applied to the frontal plane. A common challenge with learning-based methods is their reliance on large datasets. To mitigate this while leveraging the existing algorithm, we focus on fine-tuning the existing critical fault classifier, lead time classifier, and lead time regressor. Our goal is to improve performance while minimizing the need for additional training data, thus striking a balance between performance and data efficiency.
Fine-tuning is a widely used machine learning technique that adapts a pre-trained model to new data by leveraging previously learned features while adjusting to the specific requirements of the new task. This approach is both efficient and effective, especially when the new data closely resembles the original data used for pre-training. Compared to training a model from scratch, fine-tuning enables rapid adaptation to new tasks with significantly less computational effort [33].
@@ V-A Fine-Tuned Critical Fault Classifier Results
Given the nominal critical fault classifier’s poor performance on unsafe side trajectories, as discussed in Section IV-B, it is fine-tuned with 200 non-angled abrupt and 200 non-angled incipient faults, each applied to region B of side left and right, as shown in Figure 5. These regions are selected because they represent the""midpoint"" of the possible pushing areas on the sides.
Compared to the nominal critical fault classifier, the average false positive rates for the fine-tuned critical fault classifier, as shown in Table II, decreases significantly—from 0.08 to 0.0 for non-angled faults, and from 0.08 to 0.003 for angled faults. As expected, due to the direct relationship between false positive rate and lead time, the average lead time decreases from 2.54s to 1.18s and from 2.41s to 1.26s for non-angled and angled critical faults, respectively. On the other hand, the average response time increases by 0.16s and 0.17s, respectively, for non-angled and angled critical faults.
When evaluated on in-distribution data, the fine-tuned model achieves a 0.0 average false positive rate, similar to the nominal model. However, the average lead time and response time decrease and increase by 0.34s and 0.18s, respectively. Despite this, the fine-tuned critical fault classifier model achieves lead times significantly above 0.2s and nearly 0 false positive rates for all critical faults, meeting the objective outlined in [1] of detecting critical faults with sufficient lead time.
@@ V-B Fine-Tuned Lead Time Classifier Results
The lead time classifier is fine-tuned using data from the same regions as the critical fault classifier. After fine-tuning, the lead time classifier’s accuracy, as shown in Table III, increases from 0.82 to 0.96 and from 0.84 to 0.95 for the non-angled and angled critical faults in the $[0,1]$ interval, respectively. The accuracy also increases from 0.57 to 0.84 and from 0.63 to 0.80 for the $(1,2]$ interval. However, the accuracy decreases from 0.61 to 0.30 and from 0.55 to 0.27, respectively, for the $(2,H]$. This decline is expected and acceptable given that the number of data points decreases exponentially as lead time increases [1].
When evaluated on in-distribution data, the fine-tuned model’s performance slightly decreases. The accuracies drop by 0.02, 0.13, and 0.3 for the $[0,1]$, $(1,2]$, and $(2,H]$ intervals, respectively. While the drop in performance for the $(2,H]$ interval might seem significant, it is justifiable due to the previously mentioned exponential relationship between lead time and data availability. Furthermore, given the long lead times in this interval, exceeding the 0.2s threshold [1], it is reasonable to disregard the classifier’s output for this region, as it will not have severe consequences before transitioning to intervals with more accurate predictions.
@@ V-C Fine-Tuned Lead Time Regressor Results
Fine-tuning the lead time regressor using the same data as the critical fault and lead time classifier results in only acceptable performance for non-angled critical faults. Therefore, the lead time regressor is instead fine-tuned with 200 angled abrupt faults and 200 angled incipient faults, applied to the following regions: side left B and C, side right B and C, and torso front C. When evaluated on the in-distribution data, the maximum, mean, and median differences, as shown in Table IV, increase by 0.09s, 0.03s, and 0.02s, respectively, compared to the nominal model. However, when assessed on non-angled critical faults, the maximum, mean, and median differences decrease from 1.79s to 0.23s, from 0.23s to 0.03s, and from 0.07s to 0.02s, respectively, for non-angled critical faults. Similarly, for angled critical faults, the differences decrease from 2.41s to 0.32s, from 0.25s to 0.03s, and from 0.08s to 0.02s. Despite the substantial improvement in maximum difference for both angled and non-angled critical faults, the maximum difference of 0.32s achieved for angled critical faults may appear significant within the $[0,1]$ interval; however, the classifier attains this maximum difference in only 1 percent of the unsafe trajectories. This, along with the notable decrease in maximum difference, underscores the lead time regressor fine-tuned model’s capability to perform well across a wide range of critical faults.

### v3

@@ III-D3 Simulation vs Hardware — Online
IV Analysis of Nominal Fall Prediction Algorithm on Omnidirectional Faults
@@
Following the impressive online performance of the nominal fall prediction algorithm, we now evaluate its robustness to omnidirectional faults. The nominal fall prediction algorithm was trained on both abrupt and incipient faults introduced by applying pushing forces to region F at the back of the torso, as shown in Fig. 6, using features defined exclusively within the sagittal plane. As such, it is not expected to generalize well to faults outside of its training distribution, such as those introduced by pushing the robot on its sides [DBLP:journals/corr/HendrycksG16c]. However, since the training set included several trajectories in which the robot exhibited rocking motions before falling forward or backward within the sagittal plane, it remains an open question whether the algorithm can adapt to previously unseen fault conditions, such as those caused by forces applied at different locations and angles. To address this, we evaluate the performance of the three components of the fall prediction algorithm across a broad range of faults. We limit our analysis here to abrupt and incipient faults for reasons similar to those discussed in Section III. In addition, the evaluation is performed offline. However, given the negligible difference in performance between online and offline execution, as presented in Section III, we can reasonably conclude that the offline results presented here are indicative of the algorithm’s performance in real-time, online scenarios. Furthermore, the results in [ref:mungai2024fall] show that the performance of the individual components directly translates to the overall performance of the algorithm. Note that the Digit open-source dataset, [ref:digit_dataset], will be extended to include the trajectories used here.
IV-A Data Generation of Omni-directional faults
@@
To assess the algorithm’s robustness across a wider range of faults, abrupt and incipient faults are emulated using pushing forces as described in Section III. The test dataset is generated by applying both angled and non-angled forces at 16 distinct locations on the robot, distributed across the sagittal and frontal planes. These locations are grouped into four regions: torso back, torso front, right side, and left side, with five points on the back, five on the front, and three on each side. Fig. 6 illustrates the regions and points where the forces are applied.
@@
For each of the 16 locations, 100 trajectories are generated for both abrupt and incipient faults, resulting in 200 trajectories per location, using both angled and non-angled pushes. For convenience, we term critical faults that are introduced by angled and non-angled forces as critical angled and non-angled faults, respectively. Angled forces are applied by selecting the magnitudes of the x and y components of the force, as well as the angle of the force relative to the transverse plane of the robot. The x-component magnitude is chosen such that applying a non-angled x force causes a fall in approximately half of the trajectories (for both the torso back and front). Similarly, the y-component magnitude is set to produce a fall in about half of the trajectories when non-angled y forces are applied (for the torso sides). The angle of the force relative to the transverse plane is randomly selected from a uniform distribution within the range of [-45°, 45°]. These angles were determined through empirical testing, where the physical robot was pushed at various angles to identify the range within which it could be effectively disturbed. To ensure slight oscillations are present in the robot’s standing posture, random impulsive forces are introduced, but only windows during and after the fault introduction are used for fall prediction.
@@ IV-B Nominal Critical Fault Classifier Results
In [ref:mungai2024fall], the critical fault classifier demonstrated impressive results by detecting abrupt and incipient faults, achieving an average lead time of 1.52 seconds, an average response time of 0.44 seconds, and an average false positive rate of 0 when evaluated on unsafe trajectories derived in a similar manner to its training set. We refer to these results as in-distribution results.
When assessed on the broader range of abrupt and incipient critical faults, the classifier achieves an average response time comparable to the in-distribution results, as shown in Table II. While its average lead times decrease by 0.17s and 0.09s for non-angled and angled critical faults, respectively, the corresponding average false positive rates also rise to 0.08 for both cases. This indicates a decline in the nominal critical fault classifier’s performance. This decline in performance is particularly apparent for side pushes, as the false positive rates for critical faults applied to the sides of the robot are on average 0.13 for both angled and non-angled critical faults, whereas for critical faults introduced to the torso, the rates are 0.02 and 0.04 for non-angled and angled forces, respectively. These results align with expectations, as the model’s training data primarily focuses on faults introduced along the sagittal plane and does not adequately capture the variability of side or angled faults.
@@
Metrics Non Ang Ang. In Dist Non Ang Ang In Dist
FPR 0.08 0.08 0.0 0.0 0.003 0.0
Lead Time 1.35 1.43 1.52 1.18 1.26 1.18
Response Time 0.42 0.41 0.44 0.59 0.59 0.62
TABLE II: Nominal [ref:mungai2024fall] and Fine-Tuned Critical Fault Classifier results evaluated on non-angled and angled abrupt and incipient faults introduced by forces applied at various locations on the robot. Acronyms used: Non ang (non-angled), ang (angled), and in dist (in-distribution or tests from [ref:mungai2024fall]).
@@ IV-C Nominal Lead time classifier
When evaluated on in-distribution critical faults, the lead time classifier achieved accuracies of 1.0, 0.95, and 0.74 for the $[0,1]$, $(1,2]$, and $(2,H]$ lead time intervals, respectively. However, similar to the critical fault classifier, the lead time classifier performs worse when evaluated on a broader range of critical faults. For non-angled critical faults, the classifier, as shown in Table III, achieves accuracies of 0.82, 0.57, and 0.61 for the $[0,1]$, $(1,2]$, and $(2,H]$ intervals, respectively. For angled critical faults, the classifier achieves accuracies of 0.84, 0.63, and 0.55 for the same intervals. As shown in Fig. 7, this decline in performance is particularly pronounced for side critical faults. These results align with those of the critical fault classifier and are consistent with expectations, especially considering that the lead time classifier achieves its highest accuracy within the $[0,1]$ interval. Recall that, as demonstrated in [ref:mungai2024fall], the number of available data points decreases exponentially as lead time increases. The asymmetry observed in Fig. 7 between the left and right sides can be attributed to randomness in data generation, leading to varying force magnitudes even within the same range. When critical faults from one side are emulated on the other, the lead time classifier produces similar results.
@@
Intervals Non Ang Ang In Dist Non Ang Ang In Dist
$[0,1]$ 0.82 0.84 1.0 0.96 0.95 0.98
$(1,2]$ 0.57 0.63 0.95 0.84 0.80 0.82
$[2,H)$ 0.61 0.55 0.74 0.30 0.27 0.44
TABLE III: Nominal [ref:mungai2024fall] and Fine-Tuned Lead Time Classifier results evaluated on non-angled and angled abrupt and incipient faults introduced by forces applied at various locations on the robot. Acronyms used: Non ang (non-angled), ang (angled), and in dist (in-distribution or tests from [ref:mungai2024fall]).
@@ IV-D Nominal Lead time regressor
When evaluated on in-distribution critical faults within the $[0,1]$ lead time range, the lead time regressor shows minimal error, with a maximum difference of 0.09s between predicted and actual lead time, and a mean and median difference of just 0.01s [ref:mungai2024fall]. However, similar to both the critical fault classifier and the lead time classifier, the regressor performs significantly worse on non-angled and angled critical faults. For non-angled critical faults, the maximum lead time prediction error increases to 1.79s, while the mean and median errors rise to 0.23s and 0.07s, respectively, as shown in Table IV. For critical angled faults, the maximum error increases to 2.41s, with mean and median errors of 0.25s and 0.11s, respectively. Given that the regressor’s objective is to predict lead time within the $[0,1]$ interval, these error increases are unacceptable. Despite these significant errors, as shown in Fig. 8, the regressor performs better for the torso back non-angled critical faults compared to the other regions and faults.
@@
Figure 8: Lead time Regressor results when evaluated on non-angled and angled critical faults across various locations. The circles and squares indicate the non-angled and angled critical faults, respectively.
@@
Differences Non Ang Ang In Dist Non Ang Ang In Dist
Max 1.79 2.41 0.09 0.23 0.32 0.18
Mean 0.23 0.25 0.01 0.03 0.03 0.04
Median 0.07 0.11 0.01 0.02 0.02 0.03
TABLE IV: Nominal and Fine-Tuned Lead Time Regressor results evaluated on non-angled and angled abrupt and incipient faults introduced by forces applied at various locations on the robot. Acronyms used: Non ang (non-angled), ang (angled), and in dist (in-distribution or tests from [ref:mungai2024fall]).
V Robustifying Nominal Algorithm to Omnidirectional Faults
In response to the nominal fall prediction algorithm’s suboptimal performance under omnidirectional faults, particularly to faults introduced by forces applied to the frontal plane, this section explores strategies to enhance its robustness against such faults. A common challenge with learning-based methods is their reliance on large datasets. To mitigate this while leveraging the existing algorithm, we focus on fine-tuning the existing critical fault classifier, lead time classifier, and lead time regressor. Our goal is to improve performance while minimizing the need for additional training data, thus striking a balance between performance and data efficiency.
Fine-tuning is a widely used machine learning technique that adapts a pretrained model to new data by leveraging previously learned features while adjusting to the specific requirements of the new task. This approach is both efficient and effective, especially when the new data closely resembles the original data used for pretraining. Compared to training a model from scratch, fine-tuning enables rapid adaptation to new tasks with significantly less computational effort [howard2018universallanguagemodelfinetuning].
@@ V-A Fine-Tuned Critical Fault Classifier Results
Given the nominal critical fault classifier’s poor performance on unsafe side trajectories, as discussed in Section IV-B, it is fine-tuned only with 200 non-angled abrupt and 200 non-angled incipient faults, each applied to region B of side left and right, as shown in Fig. 6. These regions are selected because they represent the "midpoint" of the possible pushing areas on the sides.
Despite being fine-tuned exclusively with non-angled data from region B, the fine-tuned critical fault classifier performs significantly better than the nominal critical fault classifier across all metrics, even on unseen angled-faults. Specifically, as shown in Table II, decreases significantly - from 0.08 to 0.0 for non-angled faults, and from 0.08 to 0.003 for angled faults. As expected, due to the direct relationship between false positive rate and lead time, the average lead time decreases from 2.54s to 1.18s and from 2.41s to 1.26s for non-angled and angled critical faults, respectively. On the other hand, the average response time increases by 0.16s and 0.17s, respectively, for non-angled and angled critical faults.
When evaluated on in-distribution data, the fine-tuned model achieves a 0.0 average false positive rate, similar to the nominal model. However, the average lead time and response time decrease and increase by 0.34s and 0.18s, respectively. Despite this, the fine-tuned critical fault classifier model achieves lead times significantly above 0.2s and nearly 0 false positive rates for all critical faults, meeting the objective outlined in [ref:mungai2024fall] of detecting critical faults with sufficient lead time.
@@ V-B Fine-Tuned Lead Time Classifier Results
The lead time classifier, fine-tuned using the limited set of non-angled side data as the critical fault classifier, also shows notable improvements for all faults. As shown in Table III, its accuracy increases from 0.82 to 0.96 and from 0.84 to 0.95 for the non-angled and angled critical faults in the $[0,1]$ interval, respectively. The accuracy also increases from 0.57 to 0.84 and from 0.63 to 0.80 for the $(1,2]$ interval. However, the accuracy decreases from 0.61 to 0.30 and from 0.55 to 0.27, respectively, for the $(2,H]$. This decline is expected and acceptable given that the number of data points decreases exponentially as lead time increases [ref:mungai2024fall].
When evaluated on in-distribution data, the fine-tuned model’s performance slightly decreases. The accuracies drop by 0.02, 0.13, and 0.3 for the $[0,1]$, $(1,2]$, and $(2,H]$ intervals, respectively. While the drop in performance for the $(2,H]$ interval might seem significant, it is justifiable due to the previously mentioned exponential relationship between lead time and data availability. Furthermore, given the long lead times in this interval, exceeding the 0.2s threshold, it is reasonable to disregard the classifier’s output for this region, as it will not have severe consequences before transitioning to intervals with more accurate predictions.
@@ V-C Fine-Tuned Lead Time Regressor Results
Fine-tuning the lead time regressor using the same data as the critical fault and lead time classifier results in only acceptable performance for non-angled critical faults. Therefore, the lead time regressor is instead fine-tuned with 200 angled abrupt faults and 200 angled incipient faults, applied to the following regions: side left B and C, side right B and C, and torso front C. When evaluated on the in-distribution data, the maximum, mean, and median differences, as shown in Table IV, increase by 0.09s, 0.03s, and 0.02s, respectively, compared to the nominal model. However, when assessed on non-angled critical faults, the maximum, mean, and median differences decrease from 1.79s to 0.23s, from 0.23s to 0.03s, and from 0.07s to 0.02s, respectively, for non-angled critical faults. Similarly, for angled critical faults, the differences decrease from 2.41s to 0.32s, from 0.25s to 0.03s, and from 0.08s to 0.02s. Despite the substantial improvement in maximum difference for both angled and non-angled critical faults, the maximum difference of 0.32s achieved for angled critical faults may appear significant within the $[0,1]$ interval. However, the classifier attains this maximum difference in only 1 percent of the unsafe trajectories. This, along with the notable decrease in maximum difference, underscores the lead time regressor fine-tuned model’s capability to perform well across a wide range of critical faults.